The cost, C, in dollars, to hire landscapers to weed and seed a local park can be modelled by C equals 9 x squared minus 72 x plus 1050, where x is the number of landscapers hired to do the work. How many landscapers should be hired to minimize the cost? What is the minimum cost?
Answer by htmentor(1343)

The cost will be minimized where the derivative of the cost function is zero.
C(x) = 9x^2 - 72x + 1050
dC/dx = 0 = 18x - 72 -> x = 4
C(4) = 9*16 - 72*4 + 1050 = 906
Thus the minimum cost is $906.